(3.2.9) Listed below are the numbers of hurricanes that occurred in each year in a certain region. The data are listed in order by year. Find the​ range, variance, and standard deviation for the given sample data. Include appropriate units in the results. What important feature of the data is not revealed by any of the measures of​ variation? (include whether its in year, years^2,hurricanes,hurricane^2)
16,1,12,1,20,3,5
18,19,7,18,12,4,20
1
Expert's answer
2020-11-16T19:53:28-0500

Solution

Range:

"Range = Maximum - Minimum""= 20 -1 = 19 \space hurricanes"

Variance:

"var = {\sum {(x-\bar{x})}^2 \over n-1 }""={715.7143 \over 13} = 55. 0549 \space {hurricanes}^2"

Standard Deviation:

"sd = \sqrt{var}""=\sqrt{55. 0549} = 7.4199 \space hurricanes"

The measures of variation reveal nothing about the pattern over time